OpenStudy (anonymous):

Which is the area of a circle with a radius of 2 units? (use 3.14 for pi )

OpenStudy (anonymous):

Use the equation 2(pi)(r) which would be 2(3.14)(2) and there's your answer

OpenStudy (bookworm14):

to find the area use the formula \[\pi r^2\] your radius is 2 so do 2^2 = 4 4*3.14 = ??

OpenStudy (bookworm14):

@katiebeckett that is the formula for finding the circumference, not the area xD

OpenStudy (anonymous):

@Bookworm14 oh no oops! Sorry!

OpenStudy (bookworm14):

@katiebeckett its ok, i was helping my brother like 3 days ago and kept making the same mistake xD

OpenStudy (anonymous):

@bookworm thanks(: @katiebeckett thanks(:

OpenStudy (bookworm14):

@x_yanira your welcome!